Mylene buys a dozen cakes. the little ones cost $ 1.75 each and the kids $ 3.85 each (taxes included). Mylene paid $ 31.50 in total. How many cupcakes did Myline buy?
Answer by Boreal(15235)

x=little
12-x=kids
1.75x+3.85(12-x)=31.50
1.75x+46.20-3.85x=31.50
-2.10x=-14.70
x=7 little or $12.25
12-x=5 kids or $19.25
You can figure out which are the cupcakes.